OpenStudy (anonymous):

Find 3 consecutive odd integers such that 3 times the largest, plus the smallest, is equal to 5 less than 5 times the middle.

OpenStudy (anonymous):

x + 3(x + 4) = 5(x + 2) - 5 x + 3x + 12 = 5x + 10 - 5 4x + 12 = 5x + 5 7 = x Therefore, the smallest number is x. x = 7 7 + 2 = 9 7 + 4 = 11 Therefore, the numbers are 7, 9, and 11.

OpenStudy (anonymous):

thank youuuuu!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!